OpenStudy (anonymous):

How to solve for General Solutions ( sin2x=cosx )

hartnn (hartnn):

hint : sin 2x = 2sin x cos x.

OpenStudy (anonymous):

How did you know that from? IDs?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

wait which ID is that

hartnn (hartnn):

double angle ID

OpenStudy (anonymous):

oh rightt

OpenStudy (anonymous):

I got x=pi/2,3pi/2,5pi/6,pi/6

hartnn (hartnn):

wow, you know how to get it...thats correct but its in the range of 0 to 2pi only.

OpenStudy (anonymous):

._.

hartnn (hartnn):

general solution....means all....in terms of n....right ?

OpenStudy (anonymous):

ah yes, then you must add/subtract integer multiples of the period (2pi) to each

OpenStudy (anonymous):

alright so what would the formula be

OpenStudy (anonymous):

for the first: \[x = \frac{\pi}{2} \pm 2k\pi\] k being any integer similar for the rest